OpenStudy (marco26):

Find the equivalent D.E. by eliminating the arbitrary constant: (x-a)^2+(y-b)^2=1

OpenStudy (unklerhaukus):

unit circles at (a,b)

OpenStudy (marco26):

how?

OpenStudy (unklerhaukus):

\[(x-a)^2+(y-b)^2=1\]\[(y-b)^2=1-(x-a)^2\] \[\frac{\text d}{\text dx}\left((y-b)^2\right)=\frac{\text d}{\text dx}\left(1-(x-a)^2\right)\]\[2(y-b)\frac{\text dy}{\text dx}=-2(x-a)\]\[\frac{\text dy}{\text dx}=\frac{-2(x-a)}{2(y-b)}=\frac{a-x}{y-b}\]

OpenStudy (marco26):

yes that's the first derivative. What's next? Thanks

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx}=\frac{a-x}{y-b}\] is a first order separable differential equation

OpenStudy (unklerhaukus):

To solve the DE separate the variables \[({y-b})\cdot{\text dy}=({a-x})\cdot{\text dx}\] and integrate\[\int({y-b})\cdot{\text dy}=\int({a-x})\cdot{\text dx}\]\[\frac{y^2}2-by=ax-\frac {x^2} 2+c\]\[ y^2-2by=2ax-x^2+c\]\[(y-b)^2-b^2=-(a-x)^2+a^2+c\]\[(y-b)^2+(a-x)^2=a^2+b^2+c\] \[a^2+b^2+c=1\]